# The Impact of Responsible AI Research on Innovation and Development

### v1

@@ Introduction
Translational research plays a critical role in bridging the gap between scientific discoveries and practical applications, a process vital for technological innovation and societal advancement. This is especially pertinent in the field of Artificial Intelligence (AI), where theoretical advancements can have profound implications. For example, Vieira, O’Hagan, and O’Sullivan (2021) discussed how machine translation, a subset of AI, contributes to societal and technological innovation by applying AI in language translation for medical and legal use cases. Mayrink et al. (2022) emphasized that translational research is key to accelerating the innovation process, transforming basic science into applied science and innovation. Horowitz et al. (2017) highlighted how translational research, through “accelerators,” can rapidly generate novel solutions to eliminate disparities in healthcare, and Eweje et al. (2022) discussed its potential in biomedical research for drug discovery. In Human-Computer Interaction (HCI), Cao et al. (2023) demonstrated that HCI research significantly influences innovation, particularly systems-oriented research, through an analysis of 70 thousand patent citations. Similarly, Ahmadpoor and Jones (2017) found that fields such as nanotechnology and computer science are closely linked to patents, providing evidence of the direct impact of scientific research on technological advancement.
The research problem of studying the multi-faceted impact of Responsible AI has been long recognized. Indeed, a prominent 2022 FAccT paper called for more research on real-world impact, including the transfer of research into repositories and software products (Laufer et al. 2022). Although the effects of Responsible AI research are evident in blogs, news articles, and the drafting of forthcoming AI regulations (e.g., EU AI Act), its influence on patents and code repositories has not been fully explored. As AI becomes increasingly integrated into various sectors, it is vital to consider the ethical, fair, and sustainable development of AI technologies (Tahaei et al. 2023), and how these technologies impact the society through technological innovations and development.
To address this gap, we aim to identify the factors that contribute to the impact of Responsible AI research. [^1] Project website: https://social-dynamics.net/rai-impact Our goal is to understand two pathways (i.e., through patents and code repositories) through which Responsible AI research can drive innovation and development. In so doing, we made two main contributions:
@@
    We obtained a large dataset of 200K research papers from 1980 to 2022 in the fields of AI, Computer Vision, Data Mining, HCI, and NLP. This dataset includes 85K citations linking these papers to patents from the United States Patent and Trademark Office (USPTO), and 1 million GitHub repositories associated with these papers. Using a Sentence-Transformers Deep Learning framework for NLP and the NIST framework keywords, we identified 1,747 Responsible AI papers that served as the basis for our analysis. We developed a set of four metrics that capture the amount of research impact on patents and repositories, its time evolution, and the main factors that shape it, including authors’ institutions and papers’ conventional combination of prior knowledge.
@@
    We found that a small subset of RAI papers that go into patents or repositories is highly cited, with this process taking between 1 year for repositories up to 8 years for patents. When examining the top institutions producing impactful RAI papers in terms of patents and repositories, leading U.S. institutions are not the only contributors; significant input also comes from European and Asian institutions, reflecting a broader global influence. Finally, RAI papers typically incorporate unconventional knowledge, merging insights from diverse fields such as Machine Learning (ML) and Human-Computer Interaction (HCI).
@@
AI advances, particularly in specialized tasks such as language generation and pattern recognition, are deeply intertwined with the impact of research on innovation and development. Next, we discuss literature on Responsible AI research’s impact, and more broadly, the scientific impact on innovation and development.
@@
Abebe et al. (Abebe et al. 2020) argued that technical work can drive social change as a: 1) diagnostic tool to accurately measure social issues; 2) formalizer to define problems and explore possible solutions; 3) rebuttal to delineate the limits of what technical solutions can achieve; and 4) synecdoche by bringing public awareness to social challenges. However, as Laufer et al. (Laufer et al. 2022) caution, the journey towards leveraging Responsible AI scholarship for positive change may be hindered by economic or political interests. These interests might suppress important ethical discussions about AI (not everything that can be built ought to be), and suggest that the best solution to a problem may not necessarily be a technical one (Selbst et al. 2019) but may well require organizational cultural shift (Rakova et al. 2021; Balakrishnan et al. 2011).
@@
Responsible AI research extends its impact through various channels by raising awareness through blogs, news articles, and research papers that may not go into patents or repositories. For example, Laufer et al. conducted a meta (reflexive) study on four years of FAccT proceedings to extract research topics (e.g., group-level fairness and disinformation) and understand community’s values (e.g., industry influence over published research). Septiandri et al. (2023) analyzed 128 papers published at FAccT between 2018 and 2022, and found that 84% of these papers were exclusively based on participants from Western countries, particularly exclusively from the U.S. (63%).
@@
This points to difficulties and challenges associated with ensuring data inclusivity, mainly stemming from biases in the data collection process (Olteanu et al. 2019; Baeza-Yates 2018; Madaio et al. 2022). That is why conferences like the Neural Processing Information Systems (NeurIPS) (Ashurst et al. 2020) and the International Conference on Machine Learning (ICML) have already started mandating statements including “any risks associated with the proposed methods, methodology, application or data collection and data usage”. More recently, Olteanu et al. (2023) argued that responsible AI research needs impact statements, too; such statements aim at disclosing any possible negative consequences, contributing to more inclusive research. Additionally, researchers have proposed algorithmic impact assessments (AIAs) as a form of accountability for organizations that build and deploy automated decision-support systems (Metcalf et al. 2021), and frameworks for bridging the gap between technical and ethical aspects of AI systems (Kasirzadeh 2021).
@@
Responsible AI research draws from concepts such as responsible research and innovation. Brundage (Brundage 2016) argued that the design of AI should balance social impacts, reflect on theory, and involve public dialogue, while Owen et al. (Owen, Macnaghten, and Stilgoe 2012) emphasized innovations’ unpredictable consequences. To facilitate responsible research and innovation, Stilgoe et al. (Stilgoe, Owen, and Macnaghten 2020) proposed a framework for anticipating, reflecting, engaging, and acting upon challenges such as the narrow focus on technological innovation (Blok and Lemmens 2015). Responsible AI can significantly influence policy development. Documentation practices of datasheets for datasets (Gebru et al. 2021) and model cards (Mitchell et al. 2019) are now referenced in the draft of the EU AI Act (Act 2021), while the Ethics guidelines for Trustworthy AI (European Commission 2019) proposed by EU High-level group on Trustworthy AI based the definition of trust in scientific findings (Siau and Wang 2018). Naidoo et al. (2022) emphasized the need for a national policy framework in South Africa addressing outdated legislation, data and algorithmic bias, healthcare workforce impact, and liability dilemmas for responsible AI development and deployment. Similarly, research by Onder and Uzun (2021) demonstrated how AI could contribute to crisis management policies by aiding in various stages like preparation, mitigation-prevention, response, and recovery, providing decision support for high-quality decisions. Responsible AI research has also been used for community engagement. Verdiesen (2018) highlighted the need for guidelines and regulations in responsible AI design, development, implementation, and policy, considering its ethical, legal, and societal impact. Deshpande and Sharp (2022) identified various stakeholders in responsible AI systems, including individual, organizational, and international stakeholders, emphasizing the diverse roles and responsibilities in developing ethical AI systems.
@@
Research impact on innovation is frequently gauged through the extent to which scientific knowledge transitions into novel products and services (Jefferson et al. 2018; Cao et al. 2023; Barnes 2015; Mendoza and Sanchez 2018). Brooks (1994) emphasized the contribution of science to technology through the creation of new knowledge, tools, and techniques. The role of science in innovation has been explored by Scarrà and Piccaluga (2022), who identified several mechanisms such as technology transfer and collaboration that drive this impact. These findings illustrate the multifaceted ways in which research influences innovation, extending to fields such as nanotechnology and computer science (Ahmadpoor and Jones 2017).
A widely used source to measure research impact is through patents as they are a typical source in scholarly work to identify emerging technological innovations (Straccamore et al. 2023; Straccamore, Loreto, and Gravino 2023; Arcaute et al. 2015). Cao et al. (2023) analyzed 70,000 patent citations from leading HCI research venues, revealing a significant impact of HCI research on innovation, especially in system-oriented research, which is most often cited in patents. Tijssen (2001) investigated the influence of Dutch-authored research papers on global inventions, highlighting the interplay of knowledge transfer, including indirect research-innovation links. While their methodology is similar to our approach, they mainly used simple bibliometric counting, which might not fully reflect these indirect influences. The field of scientometrics, employing network science techniques, offers a quantitative lens to explore scientific research’s dynamics and influence (Fortunato et al. 2018; Sinha et al. 2015). By analyzing 4.8 million U.S. patents and 32 million research papers, Ahmadpoor and Jones (2017) found that mathematics was the most remote field from patents, whereas nanotechnology and computer science were among the closest. Mariani, Medo, and Lafond (2019) examined the U.S. patent citation network from 1926 to 2010 and introduced a rescaled PageRank metric, which outperformed traditional citation counts in identifying historically significant patents early. In a similar vein, Park, Leahey, and Funk (2023) studied data over six decades, encompassing 45 million papers and 3.9 million U.S. patents, and observed that recent papers and patents tend to adhere more to established science and technology trends, reducing their likelihood of deviating from traditional field norms to initiate new research directions.
@@
The impact of research on software development and code repositories, while less explored, presents an interesting dimension of technological advancement (Badashian and Stroulia 2016; Badashian et al. 2014; Blincoe et al. 2016). For example, Alqahtani, Eghan, and Rilling (2016) demonstrated how establishing traceability links in software repositories enhances knowledge sharing, highlighting the importance of cross-repository knowledge transfer for improving software security. Inokuchi et al. (2019) explored the role of academic publications as knowledge sources for software development, particularly in algorithm knowledge transfer, highlighting the direct impact of academic research on software development practices.
@@
Research Gaps. As AI continues to advance, little is known about whether the scientific efforts to ensure that AI is fair, transparent, and accountable have made a translational impact. To address this gap, this study aims to measure the impact of Responsible AI research on both innovation and development. As previously discussed, the impact may extend beyond innovation and development through blogs and news articles. However, tracking the impact on these channels requires structured data, which are currently unavailable (compared to patents and code repositories that allowed us to track the impact on innovation and development).
@@
These questions differ from traditional research questions that involve hypotheses; instead, they focus on identifying factors that contribute to the impact of RAI papers. They involve significant work in linking multiple datasets in ways that have not been done before on such a large scale. To this end, we collected, analyzed, and make publicly available [^2] Available at the project’s page: https://social-dynamics.net/rai-impact/ a large-scale dataset focusing on three elements: papers, patents, and repositories. Next, we explain the data collection and processing steps to identify RAI papers, state our method’s limitations, define the metrics of our analysis, and provide a positionality statement to clarify our perspectives on the study.
@@
Papers. From the May 2023 release of Semantic Scholar (sem 2015), we collected 200 million papers and applied a selection process to retain papers based on three criteria. The criteria for retaining papers included: a) the presence of detailed metadata fields such as ‘venue’, ‘title’, ‘year’, ‘citation count’, ‘citations’, ‘DOI’, ‘ArXiv id’, ‘authors’, ‘is open access’ and ‘abstract’; b) the publication timeframe to be between 1980 and 2022; and c) the papers to be written in English. Through this process, we compiled a dataset of 50 million papers.
Patents. From Google Patents Public Dataset, we collected 15 million patents filed in the United States Patent and Trademark Office (USPTO) between 1980 and 2022. These patents are characterized by various attributes, including the publication number, country code, date of publication, data about the inventors, the abstract, title, and references to non-patent literature. We extracted these attributes using the BigQuery service (goo 2023).
Repositories. From “Papers With Code” (pap 2023), we collected 1 million papers that had at least one link to a GitHub repository. This resulted in a set of papers from the inception of GitHub in 2008 through to 2022.
@@
We followed three steps (Figure 1). First, we curated a set of papers from a list of selected research areas and papers with code (Table B.1, Appendix). Second, we retained those papers that are about Responsible AI using a keyword-based approach based on the NIST Framework. Third, we extracted citations by matching papers referenced in patents with the Semantic Scholar metadata.
Figure 1: Overview of data processing. a) we curated papers from three sources, filtered them by year, and excluded non-research-oriented papers; b) we identified RAI papers by comparing the embeddings (generated using ST5 (Ni et al. 2021)) of paper abstracts and selected NIST keywords. c) we retrieved patent citations by matching the titles of patent references and the filtered RAI paper titles.
@@
First, we gathered an initial set of 190,790 papers. To ensure comprehensive coverage, we selected venues in which Responsible AI research is typically published (e.g., CHI, CSCW, FAccT, AIES, NeurIPS, ICML). To begin with, on Google Scholar, we selected publication venues from six core Computer Science areas: Artificial Intelligence (AI); Human-computer interaction (HCI); Natural Language Processing (NLP); Database & Information Systems; Data Mining & Analytics; and Computer Vision. For each area, we included ten highly-ranked venues based on Google Scholar’s h5 index. [^3] Google Scholar defines the h5 index as the $h$-index calculated for articles published in the most recent five complete years. Specifically, it represents the highest number $h$ such that $h$ articles from the period 2018-2022 each received at least $h$ citations. Papers from these venues cover more than 88% of the total citations in each of those six areas (Figure B.1, Appendix).
Second, we extended this initial set with all the papers (N = 950) published in the two main conferences dedicated to Responsible AI: FAccT (which started in 2018) and AIES (which started in 2017).
Third, we further added to this set with papers (N = 57,754) that had code publicly available on the site “Papers With Code”.
Fourth, we had to select the timeframe of our analysis. We determined 2015 to be an appropriate starting point. Considering the emergence of Responsible AI research and the establishment of AIES and FAccT conferences around 2017-2018, we defined a timeframe for our analysis from 2015 to 2022. This decision was based on three factors: (a) the AI community began witnessing the first significant real-world impacts of deep-learning applications around 2015 (LeCun, Bengio, and Hinton 2015; Silver et al. 2016), and the first incidents of irresponsible AI (Zhang 2015; Vincent 2016); (b) establishing a conference usually requires several years of prior planning and discussions through workshops, symposiums, and other platforms. Therefore, 2015 was a conservative choice; and (c) the entry of new technology companies such as OpenAI and Hugging Face, which possibly changed the patent landscape among technology conglomerates (e.g., IBM has shifted its focus from patent leadership in 2020 (Gil 2023)). This filtering left us with 177,881 papers.
@@
Retaining Responsible AI papers. To ensure a systematic way of retaining RAI papers in our corpus, we relied on the NIST AI Risk Management Framework (NIST 2016). The framework identifies characteristics that contribute to trustworthiness in AI systems. These characteristics are defined as a set of 31 keywords, which describe RAI topics such as fairness, explainability, accountability, privacy, and sustainability. We chose NIST because it is a recent framework from a renowned organization for developing frameworks and standards. Alternatives include the Principled Artificial Intelligence from the Berkman Klein Center (Fjeld et al. 2020), which aligns with the NIST framework.
Starting off with this list of NIST keywords, we found that some of these keywords are non-discriminative. For example, the keywords ‘validation’, ‘accuracy’, and ‘robustness’ are often used as a generic term in AI but not as a keyword that could be used to identify RAI papers (full list of removed keywords is provided in Appendix A). After removing them, we retained 12 keywords. We then added 7 more keywords related to ‘sustainability’ as it is an important topic in the public RAI debate (Strubell, Ganesh, and McCallum 2019), [^4] With the increasing power consumption needed to train AI systems, sustainability is becoming a critical focus area. This involves considering the environmental impact of AI systems, such as the energy consumption of data centers and the carbon footprint of training large models (Strubell, Ganesh, and McCallum 2019). The need for sustainable AI practices extends beyond environmental concerns to encompass long-term viability and responsible resource usage in AI development and deployment. and 8 more keywords related to governance and accountability—topics covered by regulations (e.g., EU AI Act). We then queried with the 27 keywords. During this process, we noticed spurious matches, prompting us to remove 2 keywords leading to these irrelevant matches. This process led to a final set of 25 keywords (Table A.1 in Appendix A).
@@
Retrieving citations in patents. We retrieved citations in patents in three steps. First, we connected Semantic Scholar’s paper metadata with USPTO patents. Second, we used the Grobid deep-learning framework (Grobid (gro 2011)) to extract titles and authors from patent references systematically. We then generated the embeddings of the extracted titles by using the ST5 model (Ni et al. 2021). Once we obtained the embeddings, we computed the pairwise cosine similarity to identify potential title matches, setting a distance threshold at 0.06. This threshold was chosen using the elbow rule, which involves plotting the number of title matches as a function of the distance threshold and selecting the elbow of the curve as the optimal threshold. Third, to ensure precision, we cross-referenced author names from Semantic Scholar against those in patent references using the Levenshtein distance metric (Marx and Fuegi 2022), setting a similarity threshold above 0.8 through an elbow strategy. Our approach effectively overcomes the limitations found in other methods, such as the rule-based scoring technique (rule-based scoring technique (Marx and Fuegi 2022)), which suffers from reproducibility and adaptability issues to new datasets, and a purely machine learning-based method (machine learning-based approach (bib 2022)) that is not scalable due to extensive pairwise comparisons.
Final Datasets Statistics. The final list of Responsible AI papers consists of 1,747 papers (Table 2), of which 557 (31.9%) are about fairness, 538 (30.8%) cover privacy, 318 (18.2%) discuss accountability, 219 (12.5%) focus on explainability, and 115 (6.6%) deal with sustainability.

### v4

@@ Introduction
Translational research plays a critical role in bridging the gap between scientific discoveries and practical applications, a process vital for technological innovation and societal advancement. This is especially pertinent in the field of Artificial Intelligence (AI), where theoretical advancements can have profound implications. For example, Vieira, O’Hagan, and O’Sullivan (2021) discussed how machine translation, a subset of AI, contributes to societal and technological innovation by applying AI in language translation for medical and legal use cases. Mayrink et al. (2022) emphasized that translational research is key to accelerating the innovation process, transforming basic science into applied science and innovation. Horowitz et al. (2017) highlighted how translational research, through “accelerators,” can rapidly generate novel solutions to eliminate disparities in healthcare, and Eweje et al. (2022) discussed its potential in biomedical research for drug discovery. In Human-Computer Interaction (HCI), Cao et al. (2023) demonstrated that HCI research significantly influences systems-oriented research through an analysis of 70 thousand patent citations. Similarly, Ahmadpoor and Jones (2017) found that fields such as nanotechnology and computer science are closely linked to patents, providing evidence of the direct impact of research on technological advancement.
The research problem of studying the multi-faceted impact of Responsible AI (RAI) has been long recognized. Indeed, a prominent 2022 FAccT paper called for more research on real-world impact, including the transfer of research into repositories and software products (Laufer et al. 2022). It is also equally important to recognize the translational impact of RAI research through frameworks, policies, and governance mechanisms that guide AI ethical deployment. This demonstrates the diverse pathways through which RAI research can influence societal practices and regulations (e.g., EU AI Act (Act 2021), NIST AI Risk Management Framework (NIST 2016)). Nevertheless, its influence on patents and code repositories has not been fully explored. As AI becomes increasingly integrated into various sectors, it is vital to consider the ethical, fair, and sustainable development of AI technologies (Tahaei et al. 2023), and how these technologies impact the society through technological innovations and development.
To address this gap, we aim to identify the factors that contribute to the impact of RAI. [^1] Project website: https://social-dynamics.net/rai-impact Our goal is to understand two pathways (i.e., through patents and code repositories) through which RAI research can drive innovation and development. In so doing, we made two main contributions:
@@
    We obtained a large dataset of 200K research papers from 1980 to 2022 in the fields of AI, Computer Vision, Data Mining, HCI, and NLP. This dataset includes 85K citations linking these papers to patents from the United States Patent and Trademark Office (USPTO), and 1 million GitHub repositories associated with these papers. Using a Sentence-Transformers Deep Learning framework for NLP and the NIST framework keywords, we identified 1,747 RAI papers that served as the basis for our analysis. We developed a set of four metrics that capture the amount of research impact on patents and repositories, its time evolution, and the main factors that shape it, including authors’ institutions and papers’ conventional combination of prior knowledge.
@@
    We found that a small subset of RAI papers that go into patents or repositories is highly cited, with this process taking between 1 year for repositories up to 8 years for patents. When examining the top institutions producing impactful RAI papers in terms of patents and repositories, leading U.S. institutions are not the only contributors; significant input also comes from European and Asian institutions, reflecting a broader global influence. Finally, RAI papers typically incorporate unconventional knowledge, merging insights from diverse fields such as Machine Learning (ML) and HCI.
@@
AI advances, particularly in specialized tasks such as language generation, are deeply intertwined with the impact of research on innovation and development. Next, we discuss literature on RAI research’s impact, and more broadly, the scientific impact on innovation and development.
@@
Abebe et al. (Abebe et al. 2020) argued that technical work can drive social change as a: 1) diagnostic tool to accurately measure social issues; 2) formalizer to define problems and explore possible solutions; 3) rebuttal to delineate the limits of what technical solutions can achieve; and 4) synecdoche by bringing public awareness to social challenges. However, as Laufer et al. (Laufer et al. 2022) caution, the journey towards leveraging RAI scholarship for positive change may be hindered by economic or political interests. These interests might suppress important ethical discussions (not everything that can be built ought to be), and suggest that the best solution to a problem may not necessarily be a technical one (Selbst et al. 2019) but may well require organizational cultural shift (Rakova et al. 2021; Balakrishnan et al. 2011).
@@
RAI research extends its impact through various channels by raising awareness through blogs, news articles, and research papers that may not go into patents or repositories. For example, Laufer et al. conducted a meta (reflexive) study on four years of FAccT proceedings to extract research topics (e.g., group-level fairness and disinformation) and understand community’s values (e.g., industry influence over published research). Septiandri et al. (2023) analyzed 128 papers published at FAccT between 2018 and 2022, and found that 84% of these papers were exclusively based on Western countries’ participants, particularly from the U.S. (63%).
@@
This points to difficulties and challenges associated with ensuring data inclusivity, mainly stemming from biases in the data collection process (Olteanu et al. 2019; Baeza-Yates 2018; Madaio et al. 2022). That is why conferences like the Neural Processing Information Systems (NeurIPS) (Ashurst et al. 2020) and the International Conference on Machine Learning (ICML) have already started mandating statements including “any risks associated with the proposed methods, methodology, application or data collection and data usage”. More recently, Olteanu et al. (2023) argued that RAI research needs impact statements, too; such statements aim at disclosing any possible negative consequences, contributing to more inclusive research. Additionally, researchers have proposed algorithmic impact assessments as a form of accountability for organizations that build and deploy automated decision-support systems (Metcalf et al. 2021), and frameworks for bridging the gap between technical and ethical aspects of AI systems (Kasirzadeh 2021).
@@
RAI research draws from concepts such as responsible research and innovation. Brundage (Brundage 2016) argued that the design of AI should balance social impacts, reflect on theory, and involve public dialogue, while Owen et al. (Owen, Macnaghten, and Stilgoe 2012) emphasized innovations’ unpredictable consequences. To facilitate responsible research and innovation, Stilgoe et al. (Stilgoe, Owen, and Macnaghten 2020) proposed a framework for anticipating, reflecting, engaging, and acting upon challenges such as the narrow focus on technological innovation (Blok and Lemmens 2015). RAI can significantly influence policy development. Documentation practices of datasheets for datasets (Gebru et al. 2021) and model cards (Mitchell et al. 2019) are now referenced in the draft of the EU AI Act (Act 2021), while the Ethics guidelines for Trustworthy AI (European Commission 2019) proposed by EU High-level group on Trustworthy AI based the definition of trust in scientific findings (Siau and Wang 2018). Naidoo et al. (2022) emphasized the need for a national policy framework in South Africa addressing outdated legislation, data and algorithmic bias, and liability dilemmas for responsible development and deployment. Onder and Uzun (2021) demonstrated how AI could contribute to crisis management policies by aiding in various stages like preparation, mitigation-prevention, response, and recovery, providing decision support for high-quality decisions. RAI research has also been used for community engagement. Verdiesen (2018) highlighted the need for guidelines and regulations in AI design and development, considering its ethical, legal, and societal impact. Deshpande and Sharp (2022) identified various stakeholders in AI systems, including individual, organizational, and international stakeholders, emphasizing the diverse roles and responsibilities in developing ethical AI systems.
@@
Research impact on innovation is frequently gauged through the extent to which scientific knowledge transitions into products and services (Jefferson et al. 2018; Cao et al. 2023; Barnes 2015; Mendoza and Sanchez 2018; Brooks 1994). Scarrà and Piccaluga (2022) identified several mechanisms that drive this impact such as technology transfer and collaboration, illustrating the multifaceted ways in which research influences innovation (Ahmadpoor and Jones 2017).
A widely used source to measure research impact is through patents as they are a typical source to identify emerging technological innovations (Straccamore et al. 2023; Straccamore, Loreto, and Gravino 2023; Arcaute et al. 2015). Cao et al. (2023) analyzed 70,000 patent citations from leading HCI research venues, revealing a significant impact of HCI in systems-oriented research. Tijssen (2001) investigated the influence of Dutch-authored research papers on global inventions. While their methodology is similar to our approach, they mainly used simple bibliometric counting, which might not fully reflect these indirect influences. The field of scientometrics, employing network science techniques, offers a quantitative lens to explore scientific research’s dynamics and influence (Fortunato et al. 2018; Sinha et al. 2015). By analyzing 4.8 million U.S. patents and 32 million research papers, Ahmadpoor and Jones (2017) found that mathematics was the most remote field from patents, whereas nanotechnology and computer science were among the closest. Mariani, Medo, and Lafond (2019) examined the U.S. patent citation network from 1926 to 2010 and introduced a rescaled PageRank metric, which outperformed traditional citation counts in identifying historically significant patents early. In a similar vein, Park, Leahey, and Funk (2023) studied data over six decades, encompassing 45 million papers and 3.9 million U.S. patents, and observed that recent papers and patents tend to adhere more to established science and technology trends, reducing their likelihood of deviating from traditional field norms to initiate new research directions.
@@
The impact of research on software development and code repositories, while less explored, presents an interesting dimension of technological advancement (Badashian and Stroulia 2016; Badashian et al. 2014; Blincoe et al. 2016). Alqahtani, Eghan, and Rilling (2016) demonstrated that traceability links in software repositories enhance knowledge sharing, highlighting the importance of cross-repository knowledge transfer for improving software security. Inokuchi et al. (2019) explored the role of papers as knowledge sources for software development, particularly in algorithm knowledge transfer, highlighting the direct impact of academic research on software development practices.
@@
Research Gaps. As AI continues to advance, little is known about whether the scientific efforts to ensure that AI is fair, transparent, and accountable have made a translational impact. To address this gap, this study aims to measure the impact of RAI research on both innovation and development. As previously discussed, the impact may extend beyond innovation and development through blogs and news articles. However, tracking the impact on these channels requires structured data, which are currently unavailable (compared to publicly available data on patents and code repositories).
@@
These questions differ from traditional research questions that involve hypotheses; instead, they focus on identifying factors that contribute to the impact of RAI papers. They involve significant work in linking multiple datasets in ways that have not been done before on such a large scale. To this end, we collected, analyzed, and make publicly available a large-scale dataset focusing on three elements: papers, patents, and repositories. Next, we explain the data collection and processing steps to identify RAI papers, state our method’s limitations, and define metrics for our analysis.
@@
Papers. From the May 2023 release of Semantic Scholar, [^2] https://www.semanticscholar.org/product/api we collected 200 million papers and retained papers based on three criteria: 1) the presence of detailed metadata fields such as ‘venue’, ‘title’, ‘year’, ‘citation count’, ‘citations’, ‘DOI’, ‘ArXiv id’, ‘authors’, ‘is open access’ and ‘abstract’; 2) the publication timeframe to be between 1980 and 2022; and 3) the papers to be written in English. Through this process, we compiled a dataset of 50 million papers.
Patents. From Google Patents Public Dataset, we collected 15 million patents filed in the United States Patent and Trademark Office (USPTO) between 1980 and 2022. These patents are characterized by various attributes, including the publication number, country code, date of publication, data about the inventors, the abstract, title, and references to non-patent literature. We extracted these attributes using the BigQuery service. [^3] https://console.cloud.google.com/marketplace
Repositories. From “Papers With Code”, [^4] https://paperswithcode.com/api/v1/docs we collected 1 million papers that had at least one link to a GitHub repository. This resulted in a set of papers from the inception of GitHub in 2008 through to 2022.
@@
We followed three steps (Figure 1). First, we curated a set of papers from a list of selected research areas and papers with code (Table A.5, Supplementary Information). Second, we retained those papers that are about RAI using a keyword-based approach based on the NIST Framework. Third, we extracted citations by matching papers referenced in patents with the Semantic Scholar metadata.
Figure 1: Data processing overview. a) curated papers from three sources, filtered them by year, and excluded non-research-oriented papers; b) identified RAI papers by comparing the embeddings (generated using ST5 (Ni et al. 2021)) of paper abstracts and selected NIST keywords; c) retrieved patent citations by matching the titles of patent references and those of RAI papers.
@@
First, we gathered an initial set of 190,790 papers. To ensure comprehensive coverage, we selected venues in which RAI research is typically published (e.g., CHI, CSCW, FAccT, AIES, NeurIPS, ICML). To begin with, we selected publication venues from six core Computer Science areas on Google Scholar: Artificial Intelligence (AI); Human-computer interaction (HCI); Natural Language Processing (NLP); Database & Information Systems; Data Mining & Analytics; and Computer Vision. For each area, we included ten highly-ranked venues based on Scholar’s h5 index. [^5] Google Scholar defines the h5 index as the $h$-index calculated for articles published in the most recent five complete years. Specifically, it represents the highest number $h$ such that $h$ articles from the period 2018-2022 each received at least $h$ citations. Papers from these venues cover more than 88% of the total citations in each of those six areas (Figure A.6, Supplementary Information).
Second, we extended this initial set with all the papers (N=950) published in the two main conferences dedicated to RAI: FAccT (started in 2018) and AIES (started in 2017).
Third, we added to this set papers (N=57,754) that had code publicly available on the site “Papers With Code”.
Fourth, we had to select the timeframe of our analysis. We determined 2015 to be an appropriate starting point. Considering the emergence of RAI research and the establishment of AIES and FAccT conferences around 2017-2018, we defined a timeframe for our analysis from 2015 to 2022. This decision was based on three factors: (a) the AI community began witnessing the first significant real-world impacts of deep-learning applications around 2015 (LeCun, Bengio, and Hinton 2015; Silver et al. 2016), and the first incidents of irresponsible AI (Zhang 2015; Vincent 2016); (b) establishing a conference usually requires several years of prior planning and discussions through workshops, symposiums, and other platforms. Therefore, 2015 was a conservative choice; and (c) the entry of new technology companies such as OpenAI and Hugging Face, which possibly changed the patent landscape among technology conglomerates (e.g., IBM has shifted its focus from patent leadership in 2020 (Gil 2023)). This filtering left us with 177,881 papers.
@@
Retaining RAI papers. To ensure a systematic way of retaining RAI papers in our corpus, we relied on the NIST AI Risk Management Framework (NIST 2016). The framework identifies characteristics that contribute to trustworthiness in AI systems. These characteristics are defined as a set of 31 keywords, which describe RAI topics such as fairness, explainability, accountability, privacy, and sustainability. We chose NIST because it is a recent framework from a renowned organization for developing frameworks and standards. Alternatives include the Principled Artificial Intelligence from the Berkman Klein Center (Fjeld et al. 2020), which aligns with the NIST framework.
Starting off with the NIST keywords, we found that some of these keywords are non-discriminative. For example, the keywords ‘validation’, ‘accuracy’, and ‘robustness’ are often used as a generic term in AI but not as a keyword that could be used to identify RAI papers (full list of removed keywords is provided in the Supplementary Information). After removing them, we retained 12 keywords. We then added 7 more keywords related to ‘sustainability’ as it is an important topic in the public RAI debate (Strubell, Ganesh, and McCallum 2019), [^6] With the increasing power consumption needed to train AI systems, sustainability is becoming a critical focus area. This involves considering the environmental impact of AI systems, such as the energy consumption of data centers and the carbon footprint of training large models (Strubell, Ganesh, and McCallum 2019). The need for sustainable AI practices extends beyond environmental concerns to encompass long-term viability and responsible resource usage in AI development and deployment. and 8 more keywords related to governance and accountability—topics covered by regulations (e.g., EU AI Act). We then queried with the 27 keywords. During this process, we noticed spurious matches, prompting us to remove 2 keywords leading to these irrelevant matches. This process led to a final set of 25 keywords (Table A.4, Supplementary Information).
@@
Retrieving citations in patents. We retrieved citations in patents in three steps. First, we connected Semantic Scholar’s paper metadata with USPTO patents. Second, we used the Grobid [^7] https://github.com/kermitt2/grobid deep-learning framework to extract titles and authors from patent references systematically. We then generated the embeddings of the extracted titles by using the ST5 model (Ni et al. 2021). Once we obtained the embeddings, we computed the pairwise cosine similarity to identify potential title matches, setting a distance threshold at 0.06. This threshold was chosen using the elbow rule, which involves plotting the number of title matches as a function of the distance threshold and selecting the elbow of the curve as the optimal threshold. Third, to ensure precision, we cross-referenced author names from Semantic Scholar against those in patent references using the Levenshtein distance metric (Marx and Fuegi 2022), setting a similarity threshold above 0.8 through an elbow strategy. Our approach effectively overcomes the limitations found in other methods, such as a rule-based scoring technique (Marx and Fuegi 2022) that suffers from reproducibility and adaptability issues to new datasets, or a purely machine learning-based method (Biblio Glutton) that is not scalable due to extensive pairwise comparisons.
Final Datasets Statistics. The final list of RAI papers consists of 1,747 papers (Table 2), of which 557 (31.9%) are about fairness, 538 (30.8%) cover privacy, 318 (18.2%) discuss accountability, 219 (12.5%) focus on explainability, and 115 (6.6%) deal with sustainability.